OpenStudy (anonymous):

Find the number of subsets of the set of n consecutive positive integers, {1,2,...,n} that do not contain two elements whose difference is k, where k≤n−1 This problem has been asked before and it is now closed. I will write in the comment a formula for it.

OpenStudy (anonymous):

Let f(k,n) be the number of subsets of {1,2,...,n} that do not contain two integers whose difference is k. Define the Fibonacci sequence by F(0)=F(1)=1 and F(N)=F(N-2)+F(N-1) for N=2,3,4,... The formula I get for f(k,n) is as follows. Using the division algorithm there are unique integers q,r such that n=kq+r and 0 <= r < k. Note that q=[n/k] (the greatest integer not exceeding n/k) and r=n - k*[n/k]; that is, both q and r are functions of n and k. Then f(k,n) = F(q+3)^r * F(q+2)^(k-r) One can work out a closed form for this expression using the closed form for Fibonacci numbers.

OpenStudy (anonymous):

In the case k=1, q=1, r=0 so f(1,n)= F(n+2)^2 as found by @ganeshie8

OpenStudy (anonymous):

The initial problem that was asked, k=2, n=20 so q=10, r=0 f(2,20)=F(12)^2

OpenStudy (anonymous):

@SithsAndGiggles @sidsiddhartha @Miracrown @ikram002p

OpenStudy (anonymous):

The original problem was posted by @SithsAndGiggles

ganeshie8 (ganeshie8):

Nice :) this looks more general as it avoids the dependency \(k|n\) XD

OpenStudy (sidsiddhartha):

yes thank you prof. :)

OpenStudy (anonymous):

YW @sidsiddhartha

OpenStudy (sidsiddhartha):

;)

ganeshie8 (ganeshie8):